# Complexity transitions in chaotic quantum systems: Nonstabilizerness, entanglement, and fractal dimension in SYK and random matrix models

### v2

@@ Appendix E SRE in Haar-random states
For unitary-invariant Haar-random states, the expectation value of Pauli strings can be easily calculated analytically by averaging over Haar-random states. Using the formula Roberts and Yoshida (2017)
@@
The stabilizer Rényi entropy (SRE) can be calculated using similar techniques Turkeshi et al. (2025a) and takes the following analytical forms depending on the ensemble with respect to which the Haar measure is invariant:
@@ Appendix F Pauli spectrum for RP and PLRBM bulk states
It is therefore worth investigating if relevant signatures can emerge from the Pauli spectrum, defined as the set of squared Pauli expectation values on the state Beverland et al. (2020); Turkeshi et al. (2025a):
@@
Notably, exact stabilizer states yield Pauli expectation values of either +1+1 for all $2^{N}$ stabilizers, or a balanced set of +1+1 and 1-1 values Turkeshi et al. (2025a), as can be observed in the RP bulk case (blue lines in a2). In the PLRBM model, however, bulk localized states do not align with the computational basis $\{I,Z\}^{\otimes N}$, resulting in a more dispersed Pauli spectrum (b1) compared to the RP model (a1). Even after averaging over random realizations, the PLRBM bulk state retains high variance in its Pauli spectrum across the non-ergodic and localized regimes, except in the ergodic phase where the distribution sharply concentrates Vallejo-Fabila et al. (2024).
@@ Appendix H Comparison between IPR and $D_{2}$
Although the fractal dimension $D_{2}$ can be formally related to the inverse participation ratio (IPR) $I_{2}$, defined in Eq. (1), via the expression $D_{q}=\frac{\log_{2}(I_{q})}{N(1-q)}$, relying solely on $I_{q}$ to detect phase transitions in such models can be challenging. This is because the multifractal nature of the wavefunctions manifests through the scaling relation $I_{q}\propto 1/L^{D_{q}(q-1)}$ Castellani and Peliti (1986); Mirlin et al. (1996); Mirlin and Evers (2000); Evers and Mirlin (2000); Čadež et al. (2024), which can be captured in $D_{q}$ instead of $I_{q}$. As illustrated in Fig. 11, the averaged IPR $\langle I_{2}\rangle$ reveals only the transition near $\gamma\simeq 2$, corresponding to the non-ergodic to localized transition, while it fails to resolve the ergodic to non-ergodic transition. This limitation highlights the advantage of using the fractal dimension as a more sensitive diagnostic of intermediate, multifractal phases. Moreover, taking the logarithm of the IPR is known to generally reduce the sample-to-sample fluctuations and improve the self-averaging behavior Solórzano et al. (2021).

### v3

@@ Appendix E SRE in Haar-random states
For unitary-invariant Haar-random states, the expectation value of Pauli strings can be easily calculated analytically by averaging over Haar-random states. Using the formula [164]
@@
The stabilizer Rényi entropy (SRE) can be calculated using similar techniques [77] and takes the following analytical forms depending on the ensemble with respect to which the Haar measure is invariant:
@@ Appendix F Pauli spectrum for RP and PLRBM bulk states
It is therefore worth investigating if relevant signatures can emerge from the Pauli spectrum, defined as the set of squared Pauli expectation values on the state [165, 77]:
@@
Notably, exact stabilizer states yield Pauli expectation values of either +1+1 for all $2^{N}$ stabilizers, or a balanced set of +1+1 and 1-1 values [77], as can be observed in the RP bulk case (blue lines in a2). In the PLRBM model, however, bulk localized states do not align with the computational basis $\{I,Z\}^{\otimes N}$, resulting in a more dispersed Pauli spectrum (b1) compared to the RP model (a1). Even after averaging over random realizations, the PLRBM bulk state retains high variance in its Pauli spectrum across the non-ergodic and localized regimes, except in the ergodic phase where the distribution sharply concentrates [166].
@@ Appendix H Comparison between IPR and $D_{2}$
Although the fractal dimension $D_{2}$ can be formally related to the inverse participation ratio (IPR) $I_{2}$, defined in Eq. (1), via the expression $D_{q}=\frac{\log_{2}(I_{q})}{N(1-q)}$, relying solely on $I_{q}$ to detect phase transitions in such models can be challenging. This is because the multifractal nature of the wavefunctions manifests through the scaling relation $I_{q}\propto 1/L^{D_{q}(q-1)}$ [62, 51, 57, 55, 66], which can be captured in $D_{q}$ instead of $I_{q}$. As illustrated in Fig. 11, the averaged IPR $\langle I_{2}\rangle$ reveals only the transition near $\gamma\simeq 2$, corresponding to the non-ergodic to localized transition, while it fails to resolve the ergodic to non-ergodic transition. This limitation highlights the advantage of using the fractal dimension as a more sensitive diagnostic of intermediate, multifractal phases. Moreover, taking the logarithm of the IPR is known to generally reduce the sample-to-sample fluctuations and improve the self-averaging behavior [168].